# Nonlinear Causal Discovery through a Sequential Edge Orientation Approach

### v1

@@ 11             Update $U\leftarrow E_{U}(\mathcal{G})$;
$i$$j$$k$
$i$$j$$k$
$i$$j$$k$
$i$$j$$k$(1) $k\in\text{ne}_{\mathcal{G}}(i)\cap\text{ne}_{\mathcal{G}}(j)$
$i$$j$$k$(2) $k\in\text{ne}_{\mathcal{G}}(i)\cap\text{ch}_{\mathcal{G}}(j)$
$i$$j$$k$(3) $k\in\text{ch}_{\mathcal{G}}(i)\cap\text{ne}_{\mathcal{G}}(j)$
@@
Theorem 1 shows that the edge orientation procedure in Algorithm 1 can recover the true DAG from its CPDAG. A proof is provided in the Appendix. The key of the proof is to show that there always exists an undirected edge $(i,j)$ in $\mathcal{G}$ that meets the condition in Line 1 as long as there are still undirected edges in $\mathcal{G}$. This is achieved by the careful design of the orientation rules from Line 1 to Line 1. To illustrate this point, suppose we did not include the orientation rule on Line 1 after orienting $i\rightarrow j$. As exemplified in case 1 of Figure 2, neither remaining undirected edges $i-k$ nor $j-k$ would satisfy the PANM. When evaluating $j-k$, node $i$ is now a latent parent of $k$ that would yield an error term dependent on $X_{j}$ even under the true orientation $j\rightarrow k$. The case of $i-k$ corresponds to that in Figure 1d, which does not satisfy the PANM as we discussed. Therefore, the additional orientation rules not only identify additional causal relations, but also ensure that there exists an undirected edge satisfying the PANM at the next iteration. Details on the existence of such an edge are expounded on in the proof.
@@ 3.3 Algorithm Outline
The full SNOE algorithm is formally described in Algorithm 2, which implements the key idea of Algorithm 1 through three main steps: first to learn the initial CPDAG structure, then to orient the undirected edges in the CPDAG, and lastly to remove extraneous edges. In the most general case, the final output is a PDAG. However, practitioners may choose to output a DAG if they assume it is identifiable; see Section 6.4 for further discussion.
@@
Input: Observed data $X=(X_{1},...,X_{p})$, complete undirected graph $\mathcal{G}=(V,E)$, sig. levels $\alpha_{1},\alpha_{2},\text{where }\alpha_{2}>\alpha_{1}$
@@
[…]
@@ 4       Update $E\leftarrow E\setminus\{(i,j),(j,i)\}$ and store $S_{ij}$ if found;
[…]
@@ 15 Merge edge sets to obtain $\mathcal{G}=(V,E\cup U_{\alpha_{2}})$ ;
[…]
@@ 22             Update $E\leftarrow E\setminus\{(k,i)\}$ ;
The second stage aims to determine the true causal direction of undirected edges in the CPDAG. This is accomplished through our orientation procedure OrientEdges, which finds an evaluation order for undirected edges and then identifies their causal directions. To ensure that the undirected edges are correctly oriented in a sequential manner, we develop a measure to recursively rank undirected edges accordingly to their likelihood of satisfying the independence relations implied by the PANM. Then to orient an undirected edge $X-Y$, an edge orientation test LikelihoodTest, described in Algorithm 4, computes a likelihood ratio to compare the competing directions $X\rightarrow Y$ and $Y\rightarrow X$ given their learned parent sets $\text{PA}_{X},\text{PA}_{Y}$ in the current PDAG $\mathcal{G}$. The test provides a definitive decision to either orient the edge in the preferred direction, if statistically significant, or leave it as undirected. The full details of the edge orientation procedure are presented in Algorithm 3.
[…]
@@ 4 Nonlinear Edge Orientation
To determine the true causal direction, our method employs a likelihood-ratio based test, also referred to as the edge orientation test. Given an undirected edge $X-Y$ and the parent sets $\text{PA}_{X},\text{PA}_{Y}$ in the PDAG, the test compares the bivariate conditional densities $p(X,Y\mid\text{PA}_{X},\text{PA}_{Y})$ factorized according to the directions $X\rightarrow Y$ and $Y\rightarrow X$. The likelihood test can correctly identify the causal direction when $[X,Y\mid PA_{X},PA_{Y}]$ satisfies the PANM. Furthermore, the test statistic exhibits a desirable asymptotic property that renders the results easy to obtain and interpret.
@@ 4.1 Edge Orientation Algorithm
The full edge orientation procedure is presented in Algorithm 3. Lines 3 to 3 correspond to the procedure detailed in Algorithm 1. Before ordering the edges, the algorithm partitions undirected edges $\{U_{k}\}_{k=1}^{m}$ based on the number of neighbors $|\text{ne}(U_{k})|$ shared between the two nodes on Line 3. The undirected edges are then evaluated in subsets, starting with pairs of nodes sharing $|\text{ne}(U_{k})|=0$ neighbors. Within each subset, the edges are ranked by an independence measure, which we utilize to compare their adherence to the PANM. This approach allows the algorithm to identify edges eligible for orientation more readily and reduce computation in practice, as nodes with fewer shared neighbors are more likely to satisfy the PANM. We also apply all four of Meek’s rules to further orient edges in $\mathcal{G}$, since the input PDAG may not be the true CPDAG in practice.
@@
5       Order $\widetilde{U}$ by the edge-wise independence measure (8) for each $\widetilde{U}_{i}$;
@@ 12            Apply Meek’s rules to $\mathcal{G}$ and update $U$ accordingly ;
Moreover, we utilize the undirected components of a PDAG, defined below, to facilitate parallel computation.
@@
The undirected components provide practical significance in the edge orientation procedure. They not only isolate the set of undirected edges from directed edges, but also further partition the undirected edges into disjoint sets such that undirected edges in different undirected components can be evaluated and oriented in parallel. As a result, the effects of orientations are localized, as an orientation only affects the structure of its respective undirected component.
@@ 4.2 Ranking Undirected Edges by the PANM Criterion
As demonstrated in Lemma 1, the true orientation of an undirected edge $X-Y$ is identifiable when $[X,Y\mid\text{pa}_{\mathcal{G}}(X),\text{pa}_{\mathcal{G}}(Y)]$ follows the pairwise additive noise model. Given all the undirected edges, our ranking procedure positions such edges first for orientation by utilizing an independence measure derived from the independent noise property of the PANM.
Let us employ a pairwise dependence measure $I(X,Y)$ such that $I(X,Y)=0$ if $X\perp\!\!\!\!\perp Y$ and $I(X,Y)>0$ otherwise. Let $\widehat{X}$ and $\widehat{Y}$ denote the regression estimates of $\mathbb{E}[X\mid\text{pa}_{\mathcal{G}}(X)]$ and $\mathbb{E}[Y\mid\text{pa}_{\mathcal{G}}(Y),X]$ under $X\rightarrow Y$, where $\mathcal{G}$ is a PDAG as in Algorithm 3. Then for each undirected edge $X-Y$, we calculate the maximum pairwise dependence between parents and residual of a node assuming the orientation $X\rightarrow Y$,
@@
over all $Z\in\text{pa}_{\mathcal{G}}(X)$ and $W\in\text{pa}_{\mathcal{G}}(Y)\cup\{X\}$. For the opposite orientation, $I(Y\to X)$ is calculated similarly. The maximum pairwise dependence $I(X\rightarrow Y)=0$ if the edge follows a pairwise ANM and the true orientation is $X\rightarrow Y$. Otherwise, we have $I(X\rightarrow Y)>0$. The edge-wise independence measure for $X-Y$, accounting for both possible directions, is the minimum of the two measures:
@@
Note that if edge $X-Y$ satisfies the PANM, then $\widetilde{I}(X,Y)=0$. In our work, we use normalized mutual information as the pairwise dependence measure between two random variables $Y_{1}$ and $Y_{2}$,
@@
where $MI(\cdot,\cdot)$ is the mutual information and $H(\cdot)$ is the entropy measure. This dependence measure is more comparable across different pairs of random variables, as they may have quite different entropy measures. We employ sample splitting on the data to ensure the accuracy of this metric, where the data is split into training and test sets. To calculate the quantity $I(X-\widehat{X},Z)$ in (7), for instance, we first fit a (nonlinear) regression model $\widehat{f}(\text{PA}_{X})$ for $X$ using training data. Then we obtain the fitted value $\widehat{X}=\widehat{f}(\text{PA}_{X})$ from test data. Consequently, the residual $X-\widehat{X}$ and normalized mutual information $I(X-\widehat{X},Z)$ are both calculated from test data, independent of training data, thus avoiding errors from model overfitting or reuse of the same data.
The purpose of this procedure is to distinguish edges that satisfy the PANM from those that do not. This is achieved simply by calculating $\widetilde{I}(\cdot,\cdot)$ for individual edges and sorting edges in ascending order of $\widetilde{I}(\cdot,\cdot)$. Naturally, this ranking produces an evaluation order for undirected edges, which is different from the common notion of a topological ordering of nodes. Moreover, the algorithm can orient edges that satisfy the PANM in any order, making our algorithm not order-dependent. One efficient means of evaluation is to apply Algorithm 3 separately to the undirected components $\{\tau_{1},\ldots,\tau_{T}\}$ of the input PDAG. Since they are disconnected in the PDAG, a new orientation of an edge in $\tau_{j}$ does not affect the edges in $\tau_{k}$. This also avoids iterative computation and allows for simultaneous evaluation of multiple edges. Lastly, the use of Meek’s rules in Algorithm 3 assists in reducing the number of undirected edges to sort.
@@
The edge orientation procedure is illustrated through an example in Figure 3. In the CPDAG in Figure 3b, edges $\{X_{1}-X_{2},X_{2}-X_{3},X_{4}-X_{5}\}$ follow the PANM and result in $\widetilde{I}(\cdot,\cdot)=0$. Yet since Algorithm 3 considers pairs of nodes sharing no neighbors first, it first orients $X_{1}-X_{2}$. As a result of orienting $X_{1}\rightarrow X_{2}$ and applying Meek’s rules, we obtain the maximally oriented PDAG shown in Figure 3c with several more directed edges uncovered. The algorithm would then find $\widetilde{I}(X_{3},X_{4})=0$ and $\widetilde{I}(X_{5},X_{6})=0$ because both edges satisfy the PANM, and $\widetilde{I}(X_{6},X_{7})>0$ due to $X_{5}$ missing from $\text{pa}_{\mathcal{G}}(X_{6})=\{X_{4}\}$. Therefore, our method would rank and evaluate $X_{3}-X_{4}$ and $X_{5}-X_{6}$ before $X_{6}-X_{7}$. After applying the orientation test and rules again, the algorithm would orient the last undirected edge $X_{6}-X_{7}$ to recover the true DAG, as seen in Figure 3d.
@@ 4.3 Likelihood Ratio Test for Edge Orientation
Our method adopts a comprehensive approach by considering the subgraph formed by both nodes and their learned parent sets. The likelihood ratio test returns a clear decision for edge orientation, whereas the causal relation is difficult to interpret when separate independence tests for both edge directions return statistically significant outcomes (Shah and Peters, 2020). It is also more robust against violations of the model assumptions, e.g. the noise distribution. Given that an undirected edge meets the PANM criterion, the algorithm applies the test to determine its causal direction. We first introduce the formulation of the test statistic and then describe the testing procedure.
Our orientation test takes inspiration from Vuong’s test, a series of likelihood ratio tests for model selection and testing non-nested hypotheses (Vuong, 1989). Consider two nodes $X,Y$ connected by an undirected edge in a PDAG, where $Z_{1}=\text{PA}_{X}$ and $Z_{2}=\text{PA}_{Y}$ have been identified. As indicated in Lemma 1, the PANM is only fulfilled under one causal direction. If the true direction is $X\rightarrow Y$, a reverse causal model $Y\rightarrow X$ would not satisfy the independence relations in Definition (ii) nor adequately fit to the joint distribution. Building on this insight, we compare two sets of conditional models that factorize the joint conditional density $p(x,y\mid z_{1},z_{2})$ according to the opposing directions:
Under $\text{Under }X\rightarrow Y:F_{\theta^{*}}(x,y\mid z_{1},z_{2})$ $=p(y\mid z_{2},x;\theta_{1}^{*})p(x\mid z_{1};\theta_{2}^{*}),$ (10)
Under $\text{Under }Y\rightarrow X:G_{\gamma^{*}}(x,y\mid z_{1},z_{2})$ $=p(x\mid z_{1},y;\gamma_{1}^{*})p(y\mid z_{2};\gamma_{2}^{*}).$ (11)
The two joint conditional densities are parameterized respectively by $\theta^{*}=(\theta_{1}^{*},\theta_{2}^{*})$ and $\gamma^{*}=(\gamma_{1}^{*},\gamma_{2}^{*})$. To obtain consistent estimates of the expectations of log-likelihood, $\mathbb{E}[\log F_{\theta^{*}}(X,Y\mid Z_{1},Z_{2})]$ and $\mathbb{E}[\log G_{\gamma^{*}}{(X,Y\mid Z_{1},Z_{2})}]$, we perform two-fold sample splitting on the observed data set, where the training data is used to estimate model parameters $\widehat{\theta},\widehat{\gamma}$ and the test data is used to evaluate the log-likelihood. This ensures $\widehat{\theta},\widehat{\gamma}$ are independent of the test data log-likelihood values, $\ell(\widehat{\theta}\mid X_{i},Y_{i})$ and $\ell(\widehat{\gamma}\mid X_{i},Y_{i})$, where $(X_{i},Y_{i})$ are from the test data, and allows us to establish the asymptotic distribution of the test statistic.
@@
The test uses the likelihood ratio statistic to select the model closest to the true conditional distribution. The null hypothesis $H_{0}$ indicates that the likelihood estimates should be comparable, hence we cannot identify the causal direction from the observed data. The alternative hypotheses, $H_{f}$ and $H_{g}$, are accepted when a particular edge direction is more probable. The variance of the difference in log-likelihoods computed with respect to the conditional distribution $[X,Y\mid Z_{1},Z_{2}]$ is denoted as
@@
Furthermore, let $\widehat{F}$ and $\widehat{G}$ be defined as the estimates of $F_{\theta}$ and $G_{\gamma}$ by nonparametric methods. The likelihood ratio, still denoted as $LR_{n}(\widehat{\theta},\widehat{\gamma})$ for simplicity, is calculated using nonparametric estimates $\widehat{F}$ and $\widehat{G}$ instead. Now, we establish the following properties for the likelihood ratio test:
@@
[…]
@@
[…]
@@
[…]
@@
is the smaller variance of the joint log-likelihood estimates computed under one direction. If $\widehat{\omega}_{n}^{2}/v^{2}<\delta$, for some small threshold $\delta$, then we bypass the likelihood test and declare the edge direction as indistinguishable.
Input: PDAG $\mathcal{G}$, undirected edge $X-Y$, observed data $\{X,Y,\text{PA}_{\mathcal{G}}(X),\text{PA}_{\mathcal{G}}(Y)\}$, sig. level $\alpha$

### v2

@@ 11      Update $U\leftarrow E_{U}(\mathcal{G})$;
iijjkk
iijjkk
iijjkk
iijjkk(1) $k\in\text{ne}_{\mathcal{G}}(i)\cap\text{ne}_{\mathcal{G}}(j)$
iijjkk(2) $k\in\text{ne}_{\mathcal{G}}(i)\cap\text{ch}_{\mathcal{G}}(j)$
iijjkk(3) $k\in\text{ch}_{\mathcal{G}}(i)\cap\text{ne}_{\mathcal{G}}(j)$
@@
Theorem 1 shows that the edge orientation procedure in Algorithm 1 can recover the true DAG from its CPDAG. A proof is provided in the Appendix. The key of the proof is to show that there always exists an undirected edge $(i,j)$ in $\mathcal{G}$ that meets the condition in Line 1 as long as there are still undirected edges in $\mathcal{G}$. This is achieved by the careful design of the orientation rules from Line 1 to Line 1. To illustrate this point, suppose we did not include the orientation rule on Line 1 after orienting $i\rightarrow j$. As exemplified in case 1 of Figure 2, neither remaining undirected edges iki-k nor jkj-k would satisfy the PANM. When evaluating jkj-k, node $i$ is now a latent parent of $k$ that would yield an error term dependent on $X_{j}$ even under the true orientation $j\rightarrow k$. The case of iki-k corresponds to that in Figure 1d, which does not satisfy the PANM as we discussed. Therefore, the additional orientation rules not only identify additional causal relations, but also ensure that there exists an undirected edge satisfying the PANM at the next iteration. Details on the existence of such an edge are expounded on in the proof.
@@ 3.3 Algorithm Outline
The full SNOE algorithm is formally described in Algorithm 2, which implements the key idea of Algorithm 1 through three main steps: first to learn the initial CPDAG structure, then to orient the undirected edges in the CPDAG, and lastly to remove extraneous edges. In the most general case, the final output is a PDAG. However, practitioners may choose to output a DAG if they assume it is identifiable. See Remark 2 in Section 4.3 for details.
First, we apply a modified version of the PC algorithm (spirtes1991algorithm) to learn the initial structure (Lines $\textendash2$). Specifically, we employ two significance levels: a stringent threshold $\alpha_{1}$ to learn the CPDAG and a relaxed threshold $\alpha_{2}$ to obtain a set of candidate edges. In our implementation, we use the partial correlation test to detect conditional independence relations. Starting with a complete, undirected graph, the PC algorithm removes edge $(i,j)$ if nodes $(X_{i},X_{j})$ are independent given a subset of their neighbors $S_{ij}\subseteq V$, tested at significance level $\alpha_{1}$ (Lines $\textendash 2$). To obtain the candidate edges, we in fact first learn the skeleton using the relaxed significance level $\alpha_{2}$ in the conditional independence tests, resulting in a denser skeleton as described on Lines $\textendash 2$. The candidate edges $U_{\alpha_{2}}$ (Line 2) are the edges removed when continuing the skeleton learning phase with $\alpha_{1}$, and then are reintroduced to form the graph $\mathcal{G}=(V,E\cup U_{\alpha_{2}})$. The procedure is practically equivalent to learning a CPDAG under a strict significance level, then adding undirected edges between pairs of nodes with moderate association for consideration. We essentially separate skeleton learning and edge orientation into two tasks to obtain v-structures and directed edges with higher confidence, while preserving candidate edges to reduce the number of missing edges in the graphical structure. Although our work utilizes the PC algorithm, any causal discovery algorithm that learns the equivalence class of DAG $\mathcal{G}_{0}$, with multiple sparsity levels, would be compatible with our method.
@@
3   Search for separating set $S_{ij}\subseteq V$ such that $\textit{p-val}(X_{i}\perp\!\!\!\!\perp X_{j}|S_{ij})>\alpha_{2}$;
@@ 4   Update $E\leftarrow E\setminus\{(i,j),(j,i)\}$ and store $S_{ij}$ if found;
8   Search for separating set $S_{ij}\subseteq V$ such that $\textit{p-val}(X_{i}\perp\!\!\!\!\perp X_{j}|S_{ij})>\alpha_{1}$;
@@ 15 Merge edge sets to obtain $\mathcal{G}=(V,E\cup U_{\alpha_{2}})$;
20   Construct GAM $X_{i}\sim\sum_{k\in\text{pa}_{\mathcal{G}}(i)\cup\text{ne}_{\mathcal{G}}(i)}f_{i,k}(X_{k})$ and obtain each $\textit{p-val}(f_{i,k}(X_{k}))$ from significance testing;
21 if $\textit{p-val}(f_{i,k}(X_{k}))>\alpha_{1}$ then
@@ 22      Update $E\leftarrow E\setminus\{(k,i)\}$;
The second stage aims to determine the true causal direction of undirected edges in the CPDAG. This is accomplished through our orientation procedure OrientEdges, which finds an evaluation order for undirected edges and then identifies their causal directions. To ensure that the undirected edges are correctly oriented in a sequential manner, we develop a measure to recursively rank undirected edges accordingly to their likelihood of satisfying the independence relations implied by the PANM. Then to orient an undirected edge XYX-Y, the edge orientation test LikelihoodTest, described in Algorithm 4, computes a likelihood ratio to compare the competing directions $X\rightarrow Y$ and $Y\rightarrow X$ given their learned parent sets $\text{PA}_{X},\text{PA}_{Y}$ in the current PDAG $\mathcal{G}$. The test provides a definitive decision to either orient the edge in the preferred direction, if statistically significant, or leave it as undirected. The full details of the edge orientation procedure are presented in Algorithm 3.
In the third and last step, the algorithm removes extraneous edges in the graphical structure by covariate selection (Lines $\textendash 2$). Since the graph may contain undirected edges, the algorithm also considers neighbors when performing covariate selection. Recall that a neighbor of $X$ is a node that shares an undirected edge with $X$, which excludes the parents and children of $X$ in the graph. For a node $X_{i}$, the algorithm regresses $X_{i}$ on its parents $\text{pa}_{\mathcal{G}}(X_{i})$ and neighbors $\text{ne}_{\mathcal{G}}(X_{i})$ using a generalized additive model (GAM). We perform significance testing and remove incoming edges from statistically insignificant nodes. For a neighbor $X_{j}\in\text{ne}_{\mathcal{G}}(X_{i})$, the edge is oriented as $X_{j}\rightarrow X_{i}$ if $f_{i,j}(X_{j})$ is statistically significant in the GAM for $X_{i}$ and $f_{j,i}(X_{i})$ is not significant in the GAM for $X_{j}$. If both terms are insignificant, the undirected edge is removed from the PDAG; otherwise, it remains intact.
@@ 4 Nonlinear Edge Orientation
To determine the true causal direction, our method employs a likelihood-ratio based test, also referred to as the edge orientation test. Given an undirected edge XYX-Y and the parent sets $\text{PA}_{X},\text{PA}_{Y}$ in the PDAG, the test compares the bivariate conditional densities $p(X,Y\mid\text{PA}_{X},\text{PA}_{Y})$ factorized according to the directions $X\rightarrow Y$ and $Y\rightarrow X$. The likelihood test can correctly identify the causal direction when $X,Y\mid\text{PA}_{X},\text{PA}_{Y}]$ satisfies the PANM. Furthermore, the test statistic exhibits a desirable asymptotic property that renders the result easy to obtain and interpret.
@@ 4.1 Edge Orientation Algorithm
The full edge orientation procedure is presented in Algorithm 3. Lines 3 to 3 correspond to the procedure detailed in Algorithm 1. Before ordering the edges, the algorithm partitions undirected edges $\{U_{k}\}_{k=1}^{m}$ based on the number of neighbors $|\text{ne}(U_{k})|$ shared between the two nodes on Line 3. The undirected edges are then evaluated in subsets, starting with pairs of nodes sharing $|\text{ne}(U_{k})|=0$ neighbors. Within each subset, the edges are ranked by an independence measure, which we utilize to determine their adherence to the PANM. This approach allows the algorithm to identify edges eligible for orientation more readily and reduce computation in practice, as nodes with fewer shared neighbors are more likely to satisfy the PANM. We also apply all four of Meek’s rules to further orient edges in $\mathcal{G}$, since the input PDAG may not be the true CPDAG in practice. This assists in reducing the number of undirected edges to sort.
@@
5   Order $\widetilde{U}$ by the edge-wise independence measure (8) for each $\widetilde{U}_{i}\in\widetilde{U}$;
@@ 12     Apply Meek’s rules to $\mathcal{G}$ and update $U$ accordingly;
Moreover, we may utilize the undirected components of a PDAG, defined below, to facilitate parallel orientation of undirected edges.
@@
The undirected components provide practical significance in the edge orientation procedure. They not only isolate the set of undirected edges from directed edges, but also further partition the undirected edges into disjoint sets. Since orientation of an undirected edge only affects the structure of its undirected component, edges in different undirected components can be evaluated and oriented in parallel. An efficient implementation is to apply Algorithm 3 separately to each undirected component of the input PDAG.
@@ 4.2 Ranking Undirected Edges by the PANM Criterion
As demonstrated in Lemma 1, the true orientation of an undirected edge XYX-Y is identifiable when $X,Y\mid\text{pa}_{\mathcal{G}}(X),\text{pa}_{\mathcal{G}}(Y)]$ follows the pairwise additive noise model. Given all the undirected edges, our ranking procedure positions such edges first for orientation by utilizing an independence measure derived from the independent noise property of the PANM.
Let us employ a pairwise dependence measure $I(X,Y)$ such that $I(X,Y)=0$ if $X\perp\!\!\!\!\perp Y$ and $I(X,Y)>0$ otherwise. Let $\widehat{X}$ and $\widehat{Y}$ denote the regression estimates of $\mathbb{E}[X\mid\text{pa}_{\mathcal{G}}(X)]$ and $\mathbb{E}[Y\mid\text{pa}_{\mathcal{G}}(Y),X]$ under $X\rightarrow Y$, where $\mathcal{G}$ is a PDAG as in Algorithm 3. Then for each undirected edge XYX-Y, we first calculate the maximum pairwise dependence between parents and residual of a node assuming the orientation $X\rightarrow Y$,
@@
over all $Z\in\text{pa}_{\mathcal{G}}(X)$ and $W\in\text{pa}_{\mathcal{G}}(Y)\cup\{X\}$. For the opposite orientation, $I(Y\to X)$ is calculated similarly. The maximum pairwise dependence $I(X\rightarrow Y)=0$ if the edge follows a PANM and the true orientation is $X\rightarrow Y$. Otherwise, we have $I(X\rightarrow Y)>0$. The edge-wise independence measure for XYX-Y, accounting for both possible directions, is the minimum of the two measures:
@@
Note that if edge XYX-Y satisfies the PANM, then $\widetilde{I}(X,Y)=0$. In our work, we use normalized mutual information as the pairwise dependence measure between two random variables $Y_{1}$ and $Y_{2}$,
@@
where $MI(\cdot,\cdot)$ is the mutual information and $H(\cdot)$ is the entropy measure. This dependence measure is bounded within $[0,1]$ and more comparable across different pairs of random variables, as they may have quite different or extreme entropy measures. We employ sample splitting on the data to ensure the accuracy of this metric, where the data is split into training and test sets. To calculate the quantity $I(X-\widehat{X},Z)$ in (7), for instance, we first fit a (nonlinear) regression model $\widehat{f}(\text{PA}_{X})$ for $X$ using training data. Then we obtain the fitted value $\widehat{X}=\widehat{f}(\text{PA}_{X})$ from test data. Consequently, the residual $X-\widehat{X}$ and normalized mutual information $I(X-\widehat{X},Z)$ are both calculated from test data, independent of training data, thus avoiding bias from model overfitting or reuse of the same data.
@@
The edge orientation procedure is illustrated through an example in Figure 3. In the CPDAG in Figure 3b, edges $\{X_{1}-X_{2},X_{2}-X_{3},X_{4}-X_{5}\}$ follow the PANM and result in $\widetilde{I}(\cdot,\cdot)=0$. Yet since Algorithm 3 considers pairs of nodes sharing no neighbors first, it first orients $X_{1}-X_{2}$. As a result of orienting $X_{1}\rightarrow X_{2}$ and applying Meek’s rules, we obtain the maximally oriented PDAG shown in Figure 3c with several more directed edges uncovered. Two undirected components, $\{X_{3},X_{4}\}$ and $\{X_{5},X_{6},X_{7}\}$, of the PDAG can now be oriented in parallel. The algorithm would then find $\widetilde{I}(X_{3},X_{4})=0$ and $\widetilde{I}(X_{5},X_{6})=0$ because both edges satisfy the PANM, and $\widetilde{I}(X_{6},X_{7})>0$ due to $X_{5}$ missing from $\text{pa}_{\mathcal{G}}(X_{6})=\{X_{4}\}$. Therefore, our method would rank and evaluate $X_{3}-X_{4}$ and $X_{5}-X_{6}$ before $X_{6}-X_{7}$. After applying the orientation test and rules again, the algorithm would orient the last undirected edge $X_{6}-X_{7}$ to recover the true DAG, as seen in Figure 3d.
@@ 4.3 Likelihood Ratio Test for Edge Orientation
Our method adopts a comprehensive approach by considering the subgraph formed by both nodes and their learned parent sets. The likelihood ratio test returns a clear decision for edge orientation, whereas the causal relation is difficult to interpret when separate independence tests for opposite edge directions return statistically significant outcomes (shah2020hardness). It is also more robust against violations of the model assumptions, e.g. the noise distribution. Given that an undirected edge meets the PANM criterion, the algorithm applies the test to determine its causal direction. We first introduce the formulation of the test statistic and then describe the testing procedure.
Our orientation test takes inspiration from Vuong’s test, a series of likelihood ratio tests for model selection and testing non-nested hypotheses (vuong_test). Consider two nodes $X,Y$ connected by an undirected edge in a PDAG, where $Z_{1}=\text{PA}_{X}$ and $Z_{2}=\text{PA}_{Y}$ have been identified. As indicated in Lemma 1, the PANM is only fulfilled under one causal direction. If the true direction is $X\rightarrow Y$, a reverse causal model $Y\rightarrow X$ would not satisfy the independence relations in Definition (ii) nor adequately fit to the joint distribution. Building on this insight, we compare two sets of conditional models that factorize the joint conditional density $p(x,y\mid z_{1},z_{2})$ according to the opposing directions:
$\text{Under }X\rightarrow Y:F_{\theta^{*}}(x,y\mid z_{1},z_{2})$ $=p(y\mid z_{2},x;\theta_{1}^{*})p(x\mid z_{1};\theta_{2}^{*}),$ (10)
$\text{Under }Y\rightarrow X:G_{\gamma^{*}}(x,y\mid z_{1},z_{2})$ $=p(x\mid z_{1},y;\gamma_{1}^{*})p(y\mid z_{2};\gamma_{2}^{*}).$ (11)
The two conditional densities are parameterized respectively by $\theta^{*}=(\theta_{1}^{*},\theta_{2}^{*})$ and $\gamma^{*}=(\gamma_{1}^{*},\gamma_{2}^{*})$. To obtain consistent estimates of the expectations of log-likelihood, $\mathbb{E}[\log F_{\theta^{*}}(X,Y\mid Z_{1},Z_{2})]$ and $\mathbb{E}[\log G_{\gamma^{*}}{(X,Y\mid Z_{1},Z_{2})}]$, we perform two-fold sample splitting on the observed data set, where the training data is used to estimate model parameters $\widehat{\theta},\widehat{\gamma}$ and the test data is used to evaluate the log-likelihood. This ensures that $\widehat{\theta},\widehat{\gamma}$ are independent of the test data $(X_{i},Y_{i})$ and allows us to establish the asymptotic distribution of the log-likelihoods $\ell(\widehat{\theta}\mid X_{i},Y_{i})$ and $\ell(\widehat{\gamma}\mid X_{i},Y_{i})$.
@@
The test uses the likelihood ratio statistic to select the model closest to the true conditional distribution. The null hypothesis $H_{0}$ indicates that the likelihood estimates are comparable, hence we cannot identify the causal direction from the observed data. The alternative hypotheses, $H_{f}$ and $H_{g}$, are accepted when a particular edge direction is more probable. The variance of the difference in log-likelihoods computed with respect to the conditional distribution $X,Y\mid Z_{1},Z_{2}]$ is denoted as
@@
[…]
@@
$\frac{LR_{n}(\widehat{\theta},\widehat{\gamma})}{\sqrt{n}\widehat{\omega}_{n}}\overset{D}{\longrightarrow}N(0,1),\quad\quad n\to\infty.$ (17)
@@
$\frac{LR_{n}(\widehat{\theta},\widehat{\gamma})}{\sqrt{n}\widehat{\omega}_{n}}\overset{p}{\longrightarrow}+\infty,\quad\quad n\to\infty.$ (18)
@@
$\frac{LR_{n}(\widehat{\theta},\widehat{\gamma})}{\sqrt{n}\widehat{\omega}_{n}}\overset{p}{\longrightarrow}-\infty,\quad\quad n\to\infty.$ (19)
@@
is the smaller variance of the log-likelihood estimates computed under one direction. If $\widehat{\omega}_{n}^{2}/v^{2}<\delta$, for some small threshold $\delta$, then we bypass the likelihood test and declare the edge direction as indistinguishable.
Input: PDAG $\mathcal{G}$, undirected edge XYX-Y, observed data $\{X,Y,\text{PA}_{\mathcal{G}}(X),\text{PA}_{\mathcal{G}}(Y)\}$, sig. level $\alpha$